For a four resistors n-channel JFET, find the operating points (VGS, ID, and VDS).

Assume IDSS = 5mA, VP = - 4.5V and IG ≈ 0      

Given: VDD = 12 V, R1 = 2MΩ, R2 = 1.5MΩ, RD = 6 kΩ, RSS = 4 kΩ
Answer #1

Soln 12 V Ioss = 5MA , Vp = -4,50 JIO ²6kn Io=0 2m THEO VG Vs 31.5MM 4kn YID By Voltage division VG = 1.5 mm 1.5Mn +2mn X/2 V200 ID x 4.5 a 9.64 - UKID Ich Si=a 14.142x 4.5 = 9.64-4000 x By solving x = SID 41-77x10-3 In = 1.74MA Now V5 = ID . Rs = 1.